OpenStudy (anonymous):

What is the inverse of g? g(x) = 3 x + 7

OpenStudy (anonymous):

\[\huge g(x)=\frac3{x+7}\]Is this it? ^.^

OpenStudy (anonymous):

yes @PeterPan

OpenStudy (anonymous):

Well, let's let g(x) be y, for now \[\huge y=\frac3{x+7}\]Now, what I want you to do, is switch y and x.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

\[\huge x=\frac3{y+7}\]

OpenStudy (anonymous):

And using algebra, now I want you to isolate y... What I mean is I want you to rearrange this equation, so that y is alone on one side.

OpenStudy (anonymous):

okay help

OpenStudy (anonymous):

pleae

OpenStudy (anonymous):

please

OpenStudy (anonymous):

Example?

OpenStudy (anonymous):

Say we have \[\large x=4y+16\]Then we can subtract 16 from both sides \[\large x-16 = 4y\]And then divide both sides by 4. \[\large \frac{x-16}{4}=y\]And that's how you do it ^.^